# A Survey on Adversarial Contention Resolution

### v1

@@ Abstract.
Contention resolution addresses the challenge of coordinating access by multiple processes to a shared resource such as memory, disk storage, or a communication channel. Originally spurred by challenges in database systems and bus networks, contention resolution has endured as an important abstraction for resource sharing, despite decades of technological change. Here, we survey the literature on resolving worst-case contention, where the number of processes and the time at which each process may start seeking access to the resource is dictated by an adversary. We highlight the evolution of contention resolution, where new concerns—such as security, quality of service, and energy efficiency—are motivated by modern systems. These efforts have yielded insights into the limits of randomized and deterministic approaches, as well as the impact of different model assumptions such as global clock synchronization, knowledge of the number of processors, feedback from access attempts, and attacks on the availability of the shared resource.
@@
ccs: Theory of computation Distributed algorithmsccs: Networks Network algorithmsccs: Networks Network protocolsccs: Security and privacy Network security
@@ 1. Introduction
How do we share a resource? This is a common problem, not only in large cities such as New York City, but also in a wide range of computer systems such as: wireless clients in WiFi networks (bianchi2000performance, ) and cellular networks (RamaiyanDo14, ); processes in a shared-memory architecture (Ben-DavidB17, ); nodes deciding on membership in a maximal independent set (daum2013maximal, ); and concurrency control for databases (mitra:probabilistic, ; mitra:non-exlusive, ). These are examples where system participants must share a resource, such as memory, disk storage, a data structure, or a wired or wireless communication channel. Contention resolution refers to this challenge of coordinating access to a common resource, and it arises from challenges posed by early computer systems. Over time, the problem has evolved to capture the challenges motivated by modern systems, such as wireless security, quality of service, and energy efficiency.
@@
Scope and Goals. In this survey, we focus on contention resolution in the presence of an adversary that (1) controls when participants can begin executing their contention resolution protocol, and (2) may deny access to the shared resource for periods of time; we refer to this as adversarial contention resolution. This stands in contrast to the setting where participants arrival times are governed by a stochastic process, which is an area of research that has been previously surveyed by Chlebus (chlebus2001randomized, ). The overall goal of our survey is to summarize the many models and performance metrics present in the literature on adversarial contention resolution, along with discussing their corresponding theoretical results.
@@ 1.1. Definitions, Terminology, Models, Metrics, and Algorithm Design
Several problems are synonymous with contention resolution: $k$-selection (FernandezAntaMoMu13, ), all-broadcast (chlebus2001randomized, ), the wakeup problem (jurdzinski2005probabilistic, ), and conflict resolution (komlos:asymptotically, ). To provide a unified presentation, we review terminology that is common to the literature, elaborate on performance metrics, and highlight those aspects used to organize the literature.
@@
In practice, the resource itself may be local to a single machine; for example, corresponding to the critical section of a program, or a common disk shared by multiple processes. This extends to networked settings where participants access, for instance, shared memory or separate machines via remote procedure calls.
In the literature, the resource is frequently treated as a shared communication channel, and this is the convention we use throughout this survey. Such a channel is typically referred to as a multiple access channel, which is a popular model under which several fundamental problems have been examined, such as mutual exclusion (bienkowski2010dynamic, ; czyzowicz2009consensus, ), leader election (chang:exponential-jacm, ; chang:exponential, ; nakano2002, ), consensus (czyzowicz2009consensus, ), routing (chlebus2019energy, ), broadcast problem (Marco08, ; Marco10, ; ChlebusGGPR00, ; alon1991lower, ; ClementiMS03, ; DBLP:conf/podc/Kowalski05, ; chrobak2002fast, ; piotr:explicit, ; chlebus2005almost, ; alon1991lower, ; kowalski2005time, ), and well-known random access protocols (kurose:computer, ). Here, for any given slot, accessing the resource corresponds to a participant listening (equivalently, receiving) or sending (equivalently, transmitting) a packet on the multiple access channel. Unless stated otherwise, a packet can be transmitted within a single slot. When a packet is sent in a slot, we say this packet succeeds if no other packet is sent in the same slot; otherwise, the packet fails.
For ease of exposition, we slightly abuse terminology and often refer to packets themselves taking action (e.g., sending themselves on the channel, listening on the channel), rather than introducing “participants”/“devices”/“nodes”.
@@
In contrast, in the dynamic case, packets arrive over time; this arrival model is also referred to in the literature as asynchronous (see  (DBLP:conf/podc/Kowalski05, )). The literature adopts the terminology of arrival time or activation time to indicate when a participant begins executing a contention resolution protocol in order to send packets. These times may be governed by a stochastic process (see the survey by Chlebus (chlebus2001randomized, )), or they may be scheduled by an adversary; in this survey, we focus on the latter setting.
@@
In many works, participants that listen to the channel also learn whether or not a slot contains a successful transmission. In such a setting with CD, the channel provides ternary feedback (sometimes referred to as “trinary feedback”) indicating whether the slot is empty, contains a successful transmission, or contains a collision. Similarly, in this setting, no-CD provides binary feedback. [^1] Other variants of collision detection appear in the wider literature involving multiple access channels (see  (chang:exponential, ; chang:exponential-jacm, )). Additionally, more exotic channel models have been considered; for example, De Marco and Kowalski (DBLP:conf/icdcs/MarcoK10, ) consider the case where channel feedback yields the number of senders.
There are some nuances to consider. Does a sending participant obtain different feedback with CD versus no-CD? In either case, the sender immediately learns whether its packet was sent successfully. This implies that, if the sender fails, it can infer this is the result of a collision.
Another nuance is that, while a listening participant may detect a packet success in a slot, many works do not require that the corresponding message content is accessible (see (ChangJP19, ; richa:competitive-j, ; richa:efficient-j, ; richa:jamming2, ; richa:jamming3, ; BenderFGY19, ; awerbuch:jamming, )), while for other results, this is critical (see (BenderKPY18, ; bender:contention, ; DBLP:conf/spaa/AgrawalBFGY20, )). This content is also sometimes referred to as control bits, and we refer to the property of being able to access message content snooping.
Finally, some papers consider the case where more than one channel is available to participants; this is typically referred to as a multi-channel setting. Here, participants may select any one (and only one) channel to send on or listen to in any give slot.
@@
For the contention resolution problem, the type of topology critically impacts channel feedback. With CD in a single-hop network, all participants can detect when a collision occurs. However, in a multi-hop network, only a subset of participants may detect a collision.
@@
  • Unique Identifiers. Given $n$ participants, each may have a unique identifier (ID) taken from a set of $\Theta(n)$ possible IDs. This ID may be used by an algorithm to determine when the corresponding station attempts to access the shared resource.
  • Global Clock. In the static case, all participants obtain their respective packet in the same slot $s$. This offers a common point for synchronization; that is, all participants may consider $s$ to have index 00. Thus, for the static case, there is the implicit assumption of a global clock. For dynamic arrivals, packets may be generated at participants at different times. In this case, if all participants agree on the index of the current slot, then they are said to have access to a global clock.
@@
In the static case, a natural performance measure is the number of slots until a set of packets succeeds; this number of slots is referred to as the makespan The ratio of the number of successful packets to slots is known as throughput and is another common notion of performance; note that throughput is a value between 00 and $1$. Although throughput can be differentiated from channel utilization (for example, see (throughput-vs-utilization, )), in the literature surveyed, the two are treated identically.
For the dynamic case, the above metrics remain valid when measured over periods of time during which there is at least one packet in the system However, another metric is the number of slots it takes for a participant to successfully transmit its packet starting from its activation time; this is latency. In the dynamic case, we are interested in the maximum latency over all participants which is sometimes used interchangeably with makespan in the literature. Here, we use latency to be explicit about results in the dynamic setting (specifically, in Sections 3.2 and 3.3).
@@
There is some common terminology used in the literature to characterize contention resolution algorithms, which we summarize below.
  • Deterministic versus Randomized. If the algorithm requires no random bits, then it is deterministic. Deterministic solutions are appealing from a practical standpoint in that their guarantees hold even when the number of participants and/or packets is small, admitting no error probability. In trade, they tend to provide sub-optimal asymptotic results in many settings, compared to randomized solutions.
  • Adaptive versus Non-Adaptive. An adaptive algorithm knows the full history of channel feedback and can use this information to inform future actions. Each packet contains the unique ID of the sender. Additionally, a packet may contain extra bits—referred to as control bits—that other participants may read upon receiving the packet (hradovich2020contention, ). Additionally—abusing language slightly—a packet may send messages consisting only of control bits in order to facilitate coordination (see (chlebus:adversarial, )).
@@
         In contrast, the only feedback that may alter the execution of a non-adaptive algorithm is a participant learning that its packet has been successfully transmitted, after which the participant terminates (doi:10.1137/140982763, ). There is a less-common notion of strong non-adaptiveness, where the execution of a participant is not even altered by the outcome of a transmission (again, see (doi:10.1137/140982763, )).
  • Full-sensing. Just like an adaptive algorithm, a full-sensing algorithm knows the full history of transmissions by all packets; this aligns with the notion in the literature that a full-sensing algorithm listens to the channel all the time (Goldberg-notes-2000, ). Additionally, as with an adaptive algorithm, a full-sensing algorithm may use the transmission history to inform future actions, and each packet contains a unique ID. However, a full-sensing algorithm is not allowed control bits (chlebus2006adversarial, ; hradovich2020contention, ).
  • Acknowledgement-Based. Throughout this survey, algorithms that terminate when they receive an acknowledgement are said to be acknowledgement-based (DBLP:conf/icdcs/MarcoKS19, ). In contrast, a participant may continue executing an algorithm even though its own packet has been successfully transmitted; in this case, the algorithm is non-acknowledgement-based. We note that this terminology appears to differ slightly in the literature. For example, in the stochastic dynamic model, an acknowledgement-based algorithm is one that depends on the packet’s own (rather than the full) history of transmission attempts (Goldberg-notes-2000, ). In the context of adversarial queueing theory (AQT) (see Section 3.1) have a different definition: under an acknowledgement-based algorithm, a participant sends its packet according to a schedule that can depend on its own ID and the current round (see (chlebus2006adversarial, ; DBLP:conf/icdcs/MarcoKS19, ; hradovich2020contention, )). The latter definition incorporates a dependency on IDs, since much of the work in the acknowledgement-based setting involves deterministic algorithms, which require unique IDs (see Section 3.2 for more discussion). These definitions do not contradict, but rather expand on, the definition we adopt here.
These classifications can be useful, since they indicate the strength of the underlying assumptions being made. For example, an algorithm that is randomized and full-sensing makes stronger assumptions than an algorithm that is deterministic and acknowledgement-based. A device running the former algorithm must have a pseudorandom number generator and be able to receive feedback on the channel. In contrast, the latter algorithm can function even when channel sensing is unavailable—either due to the nature of the channel or a participant’s hardware constraints—and the resulting protocol may be less complex in its implementation by eschewing randomness and knowledge of transmission history. We note that this classification can be complicated to apply in certain cases; for example, in a multi-channel setting, a device may not have the full history of feedback on all channels. In situations where this classification is clear, we present this information; otherwise, we omit it.

### v3

@@ Abstract.
Contention resolution addresses the challenge of coordinating access by multiple processes to a shared resource such as memory, disk storage, or a communication channel. Originally spurred by challenges in database systems and bus networks, contention resolution has endured as an important abstraction for resource sharing, despite decades of technological change. Here, we survey the literature on resolving worst-case contention, where the number of processes and the time at which each process may start seeking access to the resource is dictated by an adversary. We also highlight the evolution of contention resolution, where new concerns—such as security, quality of service, and energy efficiency—are motivated by modern systems. These efforts have yielded insights into the limits of randomized and deterministic approaches, as well as the impact of different model assumptions such as global clock synchronization, knowledge of the number of processors, feedback from access attempts, and attacks on the availability of the shared resource.
@@
conference: Full Conference Name; April 1, 2025; Locationbooktitle: Proceedings of the Full Conference Nameccs: Theory of computation Distributed algorithmsccs: Networks Network algorithmsccs: Networks Network protocolsccs: Security and privacy Network security
@@ 1. Introduction
How do we share a resource? This is a common problem, not only in large cities such as New York City, but also in a wide range of computer systems such as: wireless clients in WiFi networks (bianchi2000performance, ) and cellular networks (RamaiyanDo14, ); processes in a shared-memory architecture (Ben-DavidB17, ); nodes deciding on membership in a maximal independent set (daum2013maximal, ); and concurrency control for databases (mitra:probabilistic, ; mitra:non-exlusive, ). These are examples where system participants must share a resource, such as memory, disk storage, a data structure, or a wired or wireless communication channel. Contention resolution refers to this challenge of coordinating access to a common resource.
The origins of contention resolution lie with initial research on databases and old Ethernet networks (kurose:computer, ). However, the problem has grown in importance with the adoption of wireless communication, where the resource under contention is a broadcast channel. This began with the ALOHA protocol (abramson1985development, ), which initiated an effort in the research community to understand the performance of ALOHA, particularly when packet arrival times are dictated by a stochastic process (for examples, see (ALOHAnet, ; tsybakov1979ergodicity, ; ferguson1975control, ; fayolle1977stability, ; abramson1977throughput, ; rosenkrantz1983instability, ; kelly1987number, ; kelly:decentralized, )). A foundational algorithm from this time period is randomized binary exponential backoff (BEB) (MetcalfeBo76, ), which continues to play an important role in wireless communication. Subsequently, a number of variants were designed and analyzed, such as polynomial backoff (haastad1996analysis, ). The core ideas used by these older algorithms of multiplicatively increasing and decreasing the sending probability (i.e., “backing on” and “backing off”) remain key building blocks for many recent results.
While models with stochastic arrivals received significant attention from the theory community, empirical studies emerged to suggest that traffic often exhibits bursty behavior (martin2023self, ; Teymori2005, ; yu:study, ), where a large batch of packets arrive simultaneously. Thus, researchers considered a setting where a batch of packets arrive together and must be successfully transmitted on the shared channel before the next batch arrives, and so on. While this burstiness is still somewhat well-behaved—since batches never overlap—it posed new challenges over the stochastic setting. A surprising result by Bender et al. (bender:adversarial, ) demonstrated that BEB performs sub-optimally, requiring $\Omega(n\log n)$ slots to complete a single batch of $n$ packets. This spurred interest in better algorithms for batched arrivals, and we discuss these results in Section 2.
Over the past decade and a half, attention has turned to the case where the arrival times of packets are far-less constrained. One common approach has been to view this problem through the lens of adversarial queuing theory (borodin2001adversarial, ), where results are parameterized by traffic characteristics such as the arrival rate and the degree of burstiness. Another major branch in the literature addresses worst-case arrival times. Arguably, deterministic approaches have been influenced by the pioneering result by Komlos and Greenberg (komlos:asymptotically, ), leading to new solutions for worst-case arrival times. A prime example is the work by De Marco and Kowalski (doi:10.1137/140982763, ), and there has been a long line of results in this area (e.g., (DBLP:conf/icdcs/MarcoKS19, ; Chlebus:2016:SWM:2882263.2882514, ; DEMARCO20171, ; de2013contention, )). In parallel, many different randomized approaches have been explored. A recent influential paper by Chang et al. (ChangJP19, ) shows how the classic ingredients of backing off/on can yield an elegant solution to worst-case arrivals. The literature in these areas is explored in Sections 3.1, 3.2 and 3.3.
More recently, the problem has evolved further to capture the challenges motivated by modern systems. For example, energy efficiency is important to many ad-hoc and mobile network settings, where devices are typically battery-powered (SINHA201714, ; kenny:performance, ; polastre:telos, ). Another concern is malicious interference, which amounts to a denial-of-service attack (pirayesh2022jamming, ; arjoune2020smart, ). Early work in this area by Awerbuch et al. (awerbuch:jamming, ), leading to subsequent results by Richa et al. (richa:jamming2, ; DBLP:journals/dc/RichaSSZ13, ; richa:jamming3, ; DBLP:journals/ton/RichaSSZ13, ; richa:jamming4, ) and Ogierman et al. (ogierman:competitive, ; ogierman2018sade, ), showed how such interference can be tolerated, and attack-resistant contention resolution has since become an active area of research. These issues and others discussed in Sections  3.3 and 4.
@@
Scope and Goals. In this survey, we focus primarily on contention resolution in the presence of an adversary that (1) controls when participants can begin executing their contention resolution protocol, and (2) may deny access to the shared resource for periods of time; we refer to this as adversarial contention resolution. This stands in contrast to the setting where arrival times are governed by a stochastic process, which is an area of research that has been previously surveyed by Chlebus (chlebus2001randomized, ).
What motivates our focus on adversarial contention resolution? From a theoretical standpoint, it is interesting to explore what is possible in such a challenging setting, but there are also practical considerations. Network traffic does not always align with stochastic models; specifically, bursty traffic fails to be captured by well-behaved (e.g., Poisson- or Bernoulli-distributed) arrival times. Additionally, the mobility of many modern computing devices, along with connectivity challenges in wireless systems, can lead to situations where the contention for a shared resource changes rapidly over time. Such challenges are addressed by algorithms that guarantee performance results under the problem of adversarial contention resolution.
Given this motivation, our overall goal in this survey paper is to summarize the many models and performance metrics present in the literature on adversarial contention resolution, along with discussing their corresponding theoretical results.
@@ 1.1. Definitions, Terminology, Models, Metrics, and Algorithm Design
Contention resolution has gone by other names in the literature, such as $k$-selection (FernandezAntaMoMu13, ), all-broadcast (chlebus2001randomized, ), and conflict resolution (komlos:asymptotically, ). To provide a unified presentation, we review terminology that is common to the literature, elaborate on typical model assumptions and performance metrics, and highlight those aspects used to organize the literature.
@@
In practice, the resource itself may be local to a single machine; for example, corresponding to the critical section of a program, or a common disk shared by multiple processes. This extends to networked settings where participants access shared memory or separate machines via remote procedure calls.
In the literature, the resource is frequently treated as a shared communication channel, and this is the convention we use throughout this survey. Such a channel is typically referred to as a multiple access channel, which is a popular model under which several fundamental problems have been examined, such as mutual exclusion (bienkowski2010dynamic, ; czyzowicz2009consensus, ), leader election (chang:exponential-jacm, ; chang:exponential, ; nakano2002, ), consensus (czyzowicz2009consensus, ), routing (chlebus2019energy, ), broadcast problem (Marco08, ; Marco10, ; ChlebusGGPR00, ; alon1991lower, ; ClementiMS03, ; DBLP:conf/podc/Kowalski05, ; chrobak2002fast, ; piotr:explicit, ; chlebus2005almost, ; kowalski2005time, ), and well-known random access protocols (kurose:computer, ). Here, for any given slot, accessing the resource corresponds to a participant listening (equivalently, receiving) or sending (equivalently, transmitting) a packet on the multiple access channel. Unless stated otherwise, a packet can be transmitted within a single slot. When a packet is sent in a slot, we say this packet succeeds if no other packet is sent in the same slot; otherwise, the packet fails.
For ease of exposition, we will often slightly abuse terminology by speaking of packets themselves taking action (e.g., sending themselves on the channel, listening on the channel). This “packet-centric” presentation allows us to avoid referring to the corresponding “participants”/“devices”/“nodes” that actually performs these actions.
@@
In contrast, in the dynamic case, packets arrive over time; this arrival model is also referred to in the literature as asynchronous (see  (DBLP:conf/podc/Kowalski05, )). The literature adopts the terminology of arrival time or activation time to indicate when the corresponding participant begins executing a contention resolution protocol in order to send this packet. These times may be governed by a stochastic process (see the survey by Chlebus (chlebus2001randomized, )), or they may be scheduled by an adversary; in this survey, we focus on the latter setting.
@@
In many works, participants that listen to the channel also learn whether or not a slot contains a successful transmission. In such a setting with CD, the channel provides ternary feedback (sometimes referred to as “trinary feedback”) indicating whether the slot is empty, contains a successful transmission, or contains a collision. Similarly, in this setting, no-CD provides binary feedback. [^1] Other variants of CD appear in the wider literature involving multiple access channels (see  (chang:exponential, ; chang:exponential-jacm, )). Additionally, more exotic channel models have been considered; for example, De Marco and Kowalski (DBLP:conf/icdcs/MarcoK10, ) and Tsybakov (tsybakov:multiplicity, ) examine the case where channel feedback yields the number of senders.
Finally, some papers consider the case where more than one channel is available to participants; this is typically referred to as a multi-channel setting. Here, typically participants may select any one (and only one) channel to send on or listen to in any give slot. In many cases, a result involving a single channel may extend to the multi-channel setting without altering the core finding; however, there are cases where the use of multiple channels can admit significantly stronger results. Typically, this occurs when multiple channels are leveraged to achieve better latency or robustness to malicious interference, and throughout this survey, we highlight such results. In practice, multiple channels are available on several communication standards, such as IEEE 802.11 (WiFi), Bluetooth, and cellular standards (e.g., 4G and 5G LTE). Therefore, there is real-world motivation to push for results that make good use of more than one channel. However, in comparison to the single-channel setting, there are fewer results in the contention-resolution literature involving multiple channels, and so our survey is weighted more towards single-channel results.
@@
For the contention resolution problem, the type of topology may impact channel feedback. With CD in a single-hop network, all participants can detect the same channel feedback in any fixed slot. However, in a multi-hop network, it is common for different sets of participants to witness different channel feedback and, therefore, potentially update their respective states differently. Given this challenge, it is not surprising that the majority of work on contention resolution has focused the single-hop setting.
@@
  • Unique Identifiers. Given $n$ participants, each may have a unique identifier (ID) taken from a set of $\Theta(n)$ possible IDs. This ID may be used by an algorithm to determine when the corresponding station attempts to send.
  • Global Clock. In the static case, all participants obtain their respective packet in the same slot $s$. This offers a common point for synchronization; that is, all participants may consider $s$ to have index 00. Thus, for the static case, there is the implicit assumption of a global clock. For dynamic arrivals, packets may be generated at participants at different times. In this case, if all participants agree on the index of the current slot, then they are said to have access to a global clock; otherwise, they are often said to have only a local clock.
@@
In this section, we adopt the packet-centric terminology. In the static case, a natural performance measure is the number of slots until a set of packets succeeds; this number of slots is referred to as the makespan. We are typically interested in the set of all packets, but any deviations from this will be made clear from the context. For the static case, the ratio of the number of successful packets to the makespan is known as throughput, and this is another common notion of performance; note that throughput is a value between 00 and $1$.
For the dynamic case, we define a slot to be active, if at least one packet is present in the system during that slot. The above metrics remain valid when measured over active slots. However, at intermediate points in time, throughput may offer an overly-pessimistic measure of progress. For example, if a batch of $n$ packets arrive at the same time and start executing BEB, the throughput will be 00 for $\Omega(n/\log n)$ slots, despite the fact that the algorithm is progressing towards having many successful packet transmissions, which occurs after $\Theta(n)$ slots are executed.
To provide a more meaningful metric, Bender et al. (bender2020contention, ) propose the notion of “implicit throughput”. The implicit throughput at time $t$ is defined as $N_{t}/A_{t}$, where $N_{t}$ is the total number of packets that arrive at or before time $t$, and $A_{t}$ is the total number of active slots at or before time $t$. Returning to the example with a batch of $n$ packets, we note that the implicit throughput is large—specifically, $n$—in the first slot, and is $\Omega(\log n)$ over the first $\Omega(n/\log n)$ slots. The implicit throughput will continue to decrease, reaching $\Theta(1)$. Then, ultimately, the implicit throughput will be $\Theta(1/\log n)$ once all packets have succeeded, which matches the standard measure of throughput for BEB, as mentioned in Section 1. An algorithm that guarantees an implicit throughout of $\Omega(1)$ over all slots performs well, since the number of active slots so far is never asymptotically larger than the number of packets that have arrived. Furthermore, as exemplified above, at points in time where the standard notion of throughput is meaningful—such as when there are no packets in the system—the two metrics are equal.
Continuing with the dynamic case, another metric is the number of slots it takes for a participant to successfully transmit its packet starting from its activation time; this is latency. We are interested in the maximum latency over all participants. Here, we can extend throughput to be the number of packets over the maximum latency.
@@
There is some common terminology used in the literature to characterize different classes of contention resolution algorithms. An obvious division is with respect to deterministic versus randomized algorithms, where the former is characterized by the absence of random bits being used in the execution. Deterministic solutions are appealing from a practical standpoint in that their guarantees hold even when the number of participants and/or packets is small, admitting no error probability. In trade, they tend to provide sub-optimal asymptotic results in many settings, compared to randomized solutions, as we shall see.
An adaptive algorithm knows the full history of channel feedback and can use this information to inform future actions. Messages other than the sending of the packet itself are allowed; for example, such messages might pertain to estimates of the system size or coordination information. Sometimes this content is referred to as control bits (hradovich2020contention, ), and they are often used to facilitate coordination (see (chlebus:adversarial, )). In contrast, the only feedback that might alter the execution of a non-adaptive algorithm is a participant learning that its packet has been successfully transmitted, after which the participant terminates (doi:10.1137/140982763, ). There is a less-common notion of strong non-adaptiveness, where the execution of a participant is not even altered by the outcome of a transmission (again, see (doi:10.1137/140982763, )).
@@
Finer distinctions are sometimes used. A full-sensing algorithm has the same characteristics as an adaptive algorithm, but no control bits are allowed (chlebus2006adversarial, ; hradovich2020contention, ). Finally, an algorithm where packets can use acknowledgments (implicitly received after succeeding) to terminate is said to be acknowledgement-based (DBLP:conf/icdcs/MarcoKS19, ); otherwise, if packets continue executing after their respective successes, the algorithm is non-acknowledgement-based. [^2] This terminology varies slightly in the literature. In the stochastic dynamic model, an acknowledgement-based algorithm depends on the packet’s own transmission history (Goldberg-notes-2000, ). In adversarial queueing theory (AQT) (see Section 3.1), it means a packet is sent according to a schedule based on its ID and the current round (chlebus2006adversarial, ; DBLP:conf/icdcs/MarcoKS19, ; hradovich2020contention, ). The latter definition incorporates a dependency on IDs, since much of the work in the acknowledgement-based setting involves deterministic algorithms, which require unique IDs (see Section 3.2 for more discussion).
These classifications can be useful, since they indicate the strength of the underlying assumptions being made. For example, an algorithm that is randomized and full-sensing makes stronger assumptions than an algorithm that is deterministic and acknowledgement-based. A device running the former algorithm must have a pseudorandom number generator and be able to receive feedback (beyond learning of its own success) on the channel. In contrast, the latter algorithm can function even when channel sensing is unavailable—either due to the nature of the channel or a participant’s hardware constraints—and the resulting protocol may be less complex in its implementation by eschewing randomness and knowledge of transmission history.
However, we note that this classification can be complicated to apply in certain cases; for example, in a multi-channel setting, a device may not have the full history of feedback on all channels. Therefore, in situations where the classification is clear, we present this information; otherwise, we omit it.